The following reaction was carried out in a 3.75 L reaction vessel at 1100 K:

C(s)+H2O(g)⇌CO(g)+H2(g)

If during the course of the reaction, the vessel is found to contain 8.50 mol of C, 12.7 mol of H2O, 4.00 mol of CO, and 6.20 mol of H2, what is the reaction quotient Q?

Enter the reaction quotient numerically.

Q =

Consider the reaction

CO(g)+NH3(g)⇌HCONH2(g),    Kc=0.810

If a reaction vessel initially contains only CO and NH3 at concentrations of 1.00 M and 2.00 M, respectively, what will the concentration of HCONH2 be at equilibrium?

Express your answer with the appropriate units.

[HCONH2] =

For the following reaction, Kc = 255 at 1000 K.
CO (g) + Cl2 (g) ⇌ COCl2 (g)
A reaction mixture initially contains a COconcentration of 0.1550 M and a Cl2concentration of 0.177 M at 1000 K.

What is the equilibrium concentration of Cl2 at 1000 K?

Express your answer in molarity to three significant figures.

[Cl2]

Consider the following reaction:
HC2H3O2(aq)+H2O(l)⇌H3O+(aq)+C2H3O−2(aq)
Kc=1.8×10−5 at 25∘C

Part A

If a solution initially contains 0.230 M HC2H3O2, what is the equilibrium concentration of H3O+ at 25∘C?

Express your answer using two significant figures.

[H3O+] =
Answer #1

1.  reaction quotient Qc has the same expression as equilibrium constant Kc ....

Qc = [CO] [H2] / [H2O]

solids are not written in Qc expression

[CO] = no.of moles of CO/volume in litres = 4/3.75 = 1.067

[H2] = 6.2/3.75 = 1.653

[H2O] = 12.7/3.75 = 3.386

so Qc = 1.067 X 1.653 / 3.386 = 0.5209

2. Kc = [HCONH2]/[CO][NH3]
In going to equilibrium, the balanced equation says that CO and NH3 will both lose x in concentration, while HCONH2 will gain x. Substitute all we know into the equation:
0.81 = x/(1.00-x)(2.00-x),
rearranges to:
0.81x² - 3.43x + 1.62 = 0.
Using the quadratic formula, we get two answers, x = 0.541 and x = 3.69. Since [CO] cannot lose more than 1.00 M in concentration, the answer is x = [HCONH2] = 0.541M.

3.

According to the reaction equation:
CO(g) + Cl₂(g) ⇌ COCl₂(g)
the equilibrium concentrations in M are related as:
Kc = [COCl₂] / ( [CO]∙[Cl₂] )
with Kc = 255

ICE-Table
........... [CO]......... [Cl₂].......... [COCl₂]
I.......... 0.155....... 0.176............. 0
C........... -x............ -x...............+x
E....... 0.155 -x.... 0.176-x.......... x

When you substitute the expressions for the equilibrium concentrations from the last row of the table to the equilibrium equation you get:
255 = x / ( (0.155 - x)∙(0.176 - x))
<=>
255∙ (0.155 - x)∙(0.176 - x) = x
<=>
255∙(0.02728 - 0.331∙x + x²) = x
<=>
255∙x² - 85.405∙x + 6.9564 = 0
=>
x = [ 85.405 ± √( 85.405² - (4 ∙ 255 ∙ 6.9564) ) ] / ( 2 ∙ 255 )
=>
x = 0.140
or
x = 0.195
Second solution is infeasible because it would lead to negative concentrations for CO and Cl₂

Hence,  [Cl₂] = 0.176 M - x = 0.176 M - 0.140 = 0.036 M

4. So, the expression for Kc is:

Kc = 1.8 X 10^-5 = [H3O+][C2H3O2-]/[HC2H3O2]

Let x be mol/L of H3O formed. You will also form x mol/L of C2H3O2- at the same time, and you will use up x mol/L of HC2H3O2. So,

Kc = 1.8 X 10^-5 = x*x / (0.23 - x)

x = [H3O+]= 2.02*10-3 M